A square has a perimeter of 2 1/5 feet. What is the area of the square?
No idea, the fraction is throwing me off

Answer by Alan3354(69443)

A square has a perimeter of 2 1/5 feet. What is the area of the square?
--------
P = 2.2 ft
s = 2.2/4 = 0.55
Area = s^2 = 0.3025 sq ft